### v2

# MSCloudCAM: Cross-Attention with Multi-Scale Context for Multispectral Cloud Segmentation

Abstract
Clouds remain a critical challenge in optical satellite imagery, hindering reliable analysis for environmental monitoring, land cover mapping, and climate research. To overcome this, we propose MSCloudCAM, a Cross-Attention with Multi-Scale Context Network tailored for multispectral and multi-sensor cloud segmentation. Our framework exploits the spectral richness of Sentinel-2 (CloudSEN12) and Landsat-8 (L8Biome) data to classify four semantic categories: clear sky, thin cloud, thick cloud, and cloud shadow. MSCloudCAM combines a Swin Transformer backbone for hierarchical feature extraction with multi-scale context modules—ASPP and PSP—for enhanced scale-aware learning. A Cross-Attention block enables effective multisensor and multispectral feature fusion, while the integration of an Efficient Channel Attention Block (ECAB) and a Spatial Attention Module adaptively refine feature representations. Comprehensive experiments on CloudSEN12 and L8Biome demonstrate that MSCloudCAM delivers state-of-the-art segmentation accuracy, surpassing leading baseline architectures while maintaining competitive parameter efficiency and FLOPs. These results underscore the model’s effectiveness and practicality, making it well-suited for large-scale Earth observation tasks and real-world applications.
I Introduction
Clouds significantly obscure the Earth’s surface in optical satellite imagery, creating persistent challenges for remote sensing applications. Accurate detection and classification of cloud types are essential for atmospheric correction, land surface monitoring, and environmental modeling.
Current deep learning–based cloud masking and segmentation methods have demonstrated strong performance by leveraging hierarchical feature representations. However, most existing approaches remain limited in their ability to generalize across different sensors and spectral configurations. This limitation becomes especially critical in multi-class scenarios that require discriminating between thin clouds, thick clouds, and cloud shadows, where subtle spectral and spatial variations are often misclassified.
Figure 1: High-level architecture of MSCloudCAM. The model integrates a Swin Transformer backbone with ASPP and PSP for multi-scale context, and incorporates a Cross Attention module for robust feature fusion across Sentinel-2 and Landsat-8 imagery.
We propose MSCloudCAM, a novel architecture designed for robust cloud segmentation in multispectral, multi-sensor imagery. The model exploits multi-scale context encoding and cross attention to effectively fuse complementary features from different resolutions and spectral domains, enabling accurate pixel-wise classification across heterogeneous datasets.
Our key contributions are as follows:
  • We develop a hybrid backbone that integrates the Swin Transformer [1] with ASPP [2] and PSP [3] modules, enabling comprehensive multi-scale context aggregation.
  • We introduce a Cross Attention module that enhances feature fusion between high-level semantic and mid-level spatial features, improving robustness in complex cloud scenarios.
  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [4] and Landsat-8 (L8Biome) [5] imagery, consistently outperforming strong state-of-the-art segmentation networks while maintaining computational efficiency.
II Related Work
II-A Cloud Segmentation in Remote Sensing
Cloud segmentation in optical satellite imagery has long been studied due to its importance in atmospheric correction, land cover mapping, and environmental monitoring. Early rule-based methods, such as Fmask [6], applied band-specific thresholds and decision trees (e.g., cirrus, TIR bands in Landsat) but often misclassified mixed pixels, thin clouds, and bright surfaces like snow or sand [7].
Spectral index–based approaches leveraged indices such as NDSI, NDVI, and cloud probability measures [8], offering more robustness than simple thresholds but still failing under high spectral variability or semi-transparent clouds.
Later, machine learning classifiers (SVMs, Random Forests, gradient boosting) [9] combined spectral and textural features, improving accuracy but remaining limited by handcrafted inputs and shallow context modeling. These shortcomings motivated the transition to deep learning, which directly learns multi-scale, discriminative features from raw multispectral data.
II-B Deep Learning Approaches
Deep learning has become the dominant paradigm for cloud segmentation in remote sensing, offering substantial improvements in accuracy, robustness, and generalization compared to traditional threshold-based methods.
CNN-based architectures remain foundational in this domain due to their strong spatial feature extraction capabilities and hierarchical representation learning. Encoder–decoder designs such as UNet [10] and its lightweight adaptation UNetMobv2 [11] preserve fine-grained localization via skip connections, making them effective for delineating cloud boundaries. DeepLabV3+ [2] enhances context capture through atrous convolutions and the Atrous Spatial Pyramid Pooling (ASPP) module, enabling multi-scale feature aggregation for detecting clouds of varying thickness. Several CNN variants have been developed specifically for cloud detection, including CDNetv1 [12] and CDNetv2 [13] with spectral index integration and multi-branch architectures, KappaMask [14] incorporating Kappa coefficient–based loss for class imbalance handling, DBNet [15] using dual-branch feature extraction for spatial–spectral fusion, SCNN [16] employing spectral–spatial separable convolutions, MCDNet [17] with multi-channel dense connections, and HRCloudNet [18] designed for high-resolution imagery.
Transformer-based architectures leverage self-attention mechanisms to capture long-range dependencies, a property highly beneficial for clouds that span large spatial extents or exhibit non-local spectral correlations. SegFormer [19] combines the lightweight MixVision Transformer (MiT) backbone with an MLP decoder for efficient global context modeling. Mask2Former [20] unifies instance, semantic, and panoptic segmentation through masked attention, offering strong adaptability to heterogeneous cloud patterns. UPerNet-InternImage [21] integrates a hierarchical transformer backbone (InternImage) with pyramid pooling for multi-scale representation, achieving state-of-the-art results on multiple remote sensing segmentation benchmarks.
Hybrid architectures aim to combine the strengths of CNNs and transformers by using CNN layers for low-level texture and spectral edge extraction, followed by transformer modules for global reasoning. Examples include HRNet–Transformer hybrids and cross-attention fusion models [1], which have been applied to atmospheric feature segmentation.
Foundation model–based segmentation has recently emerged as a promising direction, adapting large-scale pretrained vision transformers such as DINOv2 [22] for remote sensing segmentation. These models are trained on massive, diverse datasets and provide rich, generalizable features. While they have demonstrated strong zero-shot and fine-tuned performance in RGB domains, their adaptation to multispectral or multi-sensor data requires spectral-specific preprocessing and fine-tuning strategies.
Despite these advances, most existing models are trained and evaluated on single-sensor datasets, limiting their ability to generalize across multi-sensor imagery. Moreover, few explicitly integrate multi-scale spectral–spatial features with cross-attention mechanisms tailored for cloud segmentation. Our proposed MSCloudCAM addresses both gaps by combining a Swin Transformer[1] backbone with ASPP[2] and PSP[3] modules for multi-scale context aggregation, a Cross Attention block for sensor fusion, and ECAB[23] and spatial attention[24, 25] for adaptive feature refinement.
II-C Gaps in Current Methods
Despite progress, most existing networks are trained on single-sensor datasets, limiting generalization to multi-sensor scenarios. Additionally, few works explicitly integrate multi-scale spectral-spatial features with cross-attention mechanisms for cloud segmentation. Our MSCloudCAM addresses both these gaps.
III Methodology
III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a Swin Transformer encoder with dual context modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
Figure 2: Architecture of MSCloudCAM. A multispectral Swin Transformer encoder generates hierarchical features. ASPP and PSP modules enrich deep and intermediate representations. These are fused by convolutional cross-attention, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The Swin Transformer encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
$\{f_{1},f_{2},f_{3},f_{4}\}=\mathcal{E}(\mathbf{X}),\quad f_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}},$ (1)
with channel dimensions $[96,192,384,768]$ and progressively reduced spatial sizes $(H_{i},W_{i})$. Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer detail.
III-C Multi-Scale Context Modules
To enhance representational richness, we employ two complementary context encoders that operate at different feature hierarchies.
ASPP [2]
Applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
$x_{\text{ASPP}}$ $=\phi_{\text{ASPP}}(f_{4})$
$=\bigoplus_{r\in\{1,6,12,18\}}\sigma\!\big(\text{Conv}_{3\times 3}^{(r)}(f_{4})\big)\;\oplus\;\sigma\!\big(\text{GAP}(f_{4})\big),$ (2)
where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity. This module injects large-scale semantic context while retaining spatial discrimination.
PSP [3]
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
$x_{\text{PSP}}=\phi_{\text{PSP}}(f_{3})=\bigoplus_{s\in\{1,2,3,6\}}\sigma\!\left(\text{Up}\big(\text{Pool}_{s}(f_{3})\big)\right),$ (3)
where $\text{Pool}_{s}$ denotes average pooling to an $s\times s$ grid, followed by projection and bilinear upsampling. This captures scene-level information at progressively coarser scales, which is essential for delineating fine structures such as thin clouds.
III-D Cross-Attention Fusion
The ASPP and PSP outputs are concatenated to form a joint query representation:
$x_{\text{cat}}=[\,x_{\text{ASPP}}\,\|\,x_{\text{PSP}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W}.$ (4)
A convolutional multi-head cross-attention module then refines this fused tensor by treating $x_{\text{PSP}}$ as contextual guidance:
$Q$ $=W_{Q}*x_{\text{cat}},$ (5)
$K$ $=W_{K}*x_{\text{PSP}},$ (6)
$V$ $=W_{V}*x_{\text{PSP}},$ (7)
$\text{Attn}(Q,K,V)$ $=\operatorname*{softmax}\!\Big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\Big)V,$ (8)
$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}}.$ (9)
Here, * denotes convolution, $d_{k}$ is the per-head dimension, and residual addition ensures stable learning. This mechanism aligns global semantics from ASPP with fine-grained spatial cues from PSP.
III-E Bottleneck Projection
The fused representation $\tilde{x}$ is compressed and regularized through a three-layer bottleneck:
[…]
$))),\displaystyle=\sigma\!\Big(\text{Conv}_{1\times 1}\!\big(\sigma(\text{Conv}_{3\times 3}\!(\sigma(\text{Conv}_{1\times 1}(\tilde{x}$$)\big)\Big)$, (10)
which first reduces dimensionality, then captures local spatial dependencies, and finally projects features into a compact 512-channel embedding. This step reduces redundancy and prepares the representation for attention-based refinement.
III-F Combined Attention Refinement
The bottleneck output is recalibrated by a combined attention operator:
$z^{\prime}=\phi_{\text{CA}}(z)=\phi_{\text{ECA}}(z)\odot\phi_{\text{SA}}(z)+z,$ (11)
where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [23] and $\phi_{\text{SA}}$ is spatial attention [24, 25], and $\odot$ is element-wise multiplication. This highlights discriminative regions while preserving residual information.
III-G Decoder with Deep Supervision
The refined representation $z^{\prime}$ is progressively upsampled through transposed convolutions:
$y_{i}=\sigma(\text{Deconv}_{i}(y_{i-1})),\quad i=1,\dots,4,$ (12)
with $y_{0}=z^{\prime}$. Two intermediate outputs, $\hat{Y}_{1}$ and $\hat{Y}_{2}$, are generated from $y_{1}$ and $y_{2}$, respectively, and resized to the input resolution $(H_{0},W_{0})$. The final prediction $\hat{Y}$ is obtained from $y_{4}$. The training objective jointly supervises all three outputs:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}(\hat{Y},Y)+\lambda_{1}\mathcal{L}(\hat{Y}_{1},Y)+\lambda_{2}\mathcal{L}(\hat{Y}_{2},Y),$ (13)
where $Y$ is the ground truth mask, and $(\lambda_{\text{final}},\lambda_{1},\lambda_{2})$ are supervision weights.
IV Experiments
TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [15] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
CDNetv2 [13] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
HRCloudNet [18] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
U-Net++-MobV2 90.93 86.58 49.78 66.93 73.56 95.25 92.81 66.47 80.19 83.68 94.47 96.25 95.27 96.56 95.64 91.28
DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
U-Net-ResNet101 90.62 86.16 47.15 66.05 72.50 95.08 92.56 64.08 79.56 82.82 94.24 96.13 94.97 96.57 95.48 90.95
U-Net++-ResNet101 90.90 86.43 46.24 68.46 73.01 95.23 92.72 63.24 81.27 83.12 94.45 96.16 95.06 96.74 95.60 91.20
DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
IV-A Datasets
We evaluate MSCloud on two publicly available multispectral cloud segmentation datasets.
CloudSEN12 [4] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
L8Biome [5] is derived from Landsat-8 CCA, containing 11 bands (30 m) with manual masks. The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$. The fill class is ignored. Bands are normalized with the same $3000.0$ divisor for consistency.
TABLE II: Dataset split statistics for CloudSEN12 and L8Biome.
Dataset Train Validation Test
CloudSEN12 L1C 8,500 500 1,000
CloudSEN12 L2A 8,500 500 1,000
L8Biome 13,378 4,459 4,459
IV-B Loss Function
The training objective is a weighted sum of the losses from the final prediction and the two auxiliary predictions:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}_{\text{final}}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2}.$ (14)
We set $\lambda_{\text{final}}=1.0$ and $\lambda_{1}=\lambda_{2}=0.4$, ensuring the final prediction dominates while auxiliary supervision stabilizes early-stage training.
IV-C Training Protocol
Our model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. The batch size is set to 8, and we train for 100 epochs.
IV-D Evaluation Metrics
Per-class Metrics.
$\mathrm{IoU}_{i}$ $=\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (15)
$\mathrm{F1}_{i}$ $=\frac{2\,\mathrm{TP}_{i}}{2\,\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (16)
$\mathrm{Acc}_{i}$ $=\frac{\mathrm{TP}_{i}+\mathrm{TN}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}+\mathrm{TN}_{i}}.$ (17)
Per-class metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
Ignored labels.
Pixels with sentinel values (e.g., Fill in L8Biome) are removed from $\mathbf{C}$ and all denominators via an ignore_index mask, ensuring metrics are based on identical valid-pixel sets.
Reporting.
We report class-wise IoU and F1 (clear, thick cloud, thin cloud, cloud shadow), together with mIoU, mDice, mAcc, and aAcc. Unless noted, averages are macro over classes.
V Results
V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L2A (BOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [15] 85.42 42.76 80.80 53.62 65.65 92.14 59.90 89.38 69.81 77.81 94.83 53.53 90.27 63.45 75.52 86.82
CDNetv2 [13] 85.35 44.01 80.85 54.00 66.05 92.09 61.12 89.41 70.13 78.19 95.65 56.37 88.06 63.56 75.91 86.88
HRCloudNet [18] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
U-Net++-MobV2 87.45 83.33 50.70 57.62 69.78 93.30 90.91 67.29 73.11 81.15 92.52 95.47 93.26 95.23 94.12 88.23
DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
U-Net-ResNet101 86.51 81.76 37.12 58.89 66.07 92.77 89.97 54.14 74.12 77.75 91.83 94.73 92.76 95.13 93.61 87.23
U-Net++-ResNet101 87.85 81.03 47.47 59.12 68.87 93.53 89.52 64.38 74.31 80.44 92.84 94.47 93.39 95.37 94.02 88.04
DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
Mask2Former-DINOv2 83.53 79.32 45.51 56.50 66.21 91.03 88.47 62.55 72.21 78.56 90.30 94.23 91.17 94.93 92.66 85.31
Mask2Former-Swin-T 88.91 84.35 50.72 63.25 71.81 94.13 91.51 67.30 77.49 82.61 93.61 95.77 93.33 95.30 94.50 89.01
SegFormer-MiT-B5 89.28 83.50 51.87 64.19 72.21 94.34 91.01 68.31 78.19 82.96 93.76 95.36 93.86 95.90 94.72 89.44
MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
TABLE IV: Performance comparison on L8Biome.
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
CDNetv2 [13] 73.81 24.85 75.85 00.00 43.63 84.93 39.81 86.27 00.00 52.75 90.18 31.95 89.81 00.00 52.98 78.56
HRCloudNet [18] 72.41 30.89 70.76 00.00 43.51 83.99 47.20 82.88 00.00 53.52 85.50 43.70 85.88 00.00 53.77 77.04
DBNet [15] 80.19 38.18 82.37 04.90 51.41 89.00 55.25 90.34 09.34 60.99 96.51 49.34 87.89 05.03 59.70 83.62
U-Net-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
U-Net++-MobV2 88.96 81.68 39.44 4.42 53.62 94.16 89.91 56.57 8.47 62.28 94.13 93.55 88.76 98.57 93.55 87.09
DeepLabV3+-MobV2 87.25 79.21 33.37 2.54 50.59 93.19 88.40 50.05 4.96 59.15 93.32 91.25 87.50 98.54 92.65 85.31
U-Net-ResNet101 84.29 77.00 30.56 1.25 48.28 91.47 87.01 46.81 2.48 56.94 91.30 90.94 86.06 98.01 91.58 83.16
U-Net++-ResNet101 88.16 80.23 34.94 3.81 51.78 93.71 89.03 51.78 7.34 60.46 93.73 91.94 87.87 98.52 93.01 86.03
DeepLabV3+-ResNet101 87.56 74.19 34.24 4.88 50.22 93.37 85.18 51.01 9.31 59.72 93.47 89.56 85.51 98.56 91.78 83.55
UPerNet-InternImage 54.68 46.97 18.65 1.56 30.46 60.19 53.40 26.72 2.55 35.72 91.00 90.58 83.73 98.35 90.92 81.83
Mask2Former-DINOv2 69.67 70.75 34.67 13.13 47.06 82.12 82.87 51.49 23.22 59.93 84.03 86.86 84.39 98.65 88.48 76.96
Mask2Former-Swin-T 84.36 78.64 37.50 13.54 53.51 91.52 88.04 54.55 23.85 64.49 91.74 91.62 86.01 98.62 92.00 84.00
SegFormer-MiT-B5 88.62 74.38 39.52 15.27 54.45 93.97 85.31 56.65 26.50 65.61 94.04 90.12 86.05 98.70 92.23 84.45
MSCloudCAM 88.27 82.23 40.19 21.48 58.04 93.77 90.25 57.34 35.36 69.18 93.84 92.91 88.74 98.70 93.55 87.09
V-B Qualitative Results
To further illustrate the effectiveness of the proposed method, we present visual comparisons of cloud segmentation results in Fig. 3. Each row corresponds to a different dataset, and each column corresponds to a different segmentation method. As shown in the figure, the proposed method produces more accurate delineation of thin clouds and cloud shadows, while reducing false detections compared to the baseline approaches.
Figure 3: Qualitative cloud segmentation results. Each row corresponds to a dataset. The proposed method produces sharper delineation of thin clouds and shadows compared to other approaches. Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
V-C Model Complexity Analysis
Table V summarizes the computational complexity of all evaluated models in terms of FLOPs and parameter counts. While some architectures (e.g., UPerNet-InternImage and Mask2Former-DINOv2) exhibit significantly higher complexity, our proposed MSCloudCAM achieves a favorable trade-off between accuracy and efficiency. Specifically, MSCloudCAM attains competitive performance while requiring fewer FLOPs than SegFormer-MiT-B5, and substantially fewer FLOPs than UPerNet-InternImage, with a moderate parameter count.
TABLE V: Model complexity comparison in terms of FLOPs and parameters.
Model FLOPs (G) Parameters (M)
U-Net-MobileNetV2 13.613 6.629
U-Net++-MobileNetV2 18.023 6.825
DeepLabV3+-MobileNetV2 6.057 4.379
U-Net-ResNet101 62.258 51.514
U-Net++-ResNet101 250.000 67.978
DeepLabV3+-ResNet101 56.192 45.670
UPerNet-InternImage-B-DCNv4 200–250 100–130
Mask2Former-Swin-T 73.090 47.400
Mask2Former-DINOv2-ViT-B/14 75–95 87–92
SegFormer-MiT-B5 55.430 45.652
MSCloudCAM 38.98 47.44
VI Code and Model Availability
The full implementation of our proposed MS-CloudCAM framework is publicly available. The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam. Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating transformer-based backbones with multi-scale context and attention mechanisms, our method achieves superior segmentation accuracy over strong baselines on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.
Acknowledgment
This material is based in part upon work supported by the National Science Foundation under Grants CNS-2018611 and CNS-1920182.