### v3

# MSCloudCAM: Multi-Scale Context Adaptation with Convolutional Cross-Attention for Multispectral Cloud Segmentation

Abstract
Clouds remain a major obstacle in optical satellite imaging, limiting accurate environmental and climate analysis. To address the strong spectral variability and the large scale differences among cloud types, we propose MSCloudCAM, a novel multi-scale context adapter network with convolution based cross-attention tailored for multispectral and multi-sensor cloud segmentation. A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors. And also, rather than simply stacking or concatenating their outputs, our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection. Building on this idea, we design a new convolution-based cross attention adapter that effectively fuses localized, detailed information with broader multi-scale context. Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination. Experiments on various multisensor datatsets e.g. CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8) show that MSCloudCAM outperforms recent state-of-the-art models while maintaining competitive model complexity, highlighting the novelty and effectiveness of the proposed design for large-scale Earth observation.
I Introduction
Clouds significantly obscure the Earth’s surface in optical satellite imagery, creating persistent challenges for remote sensing applications. Accurate detection and classification of cloud types are essential for atmospheric correction, land surface monitoring, and environmental modeling.
Recent deep learning–based cloud masking and segmentation methods have shown strong performance by leveraging hierarchical feature representations. However, most existing approaches struggle to generalize across heterogeneous sensors and spectral configurations. This limitation becomes especially critical in multi-class cloud segmentation requires modeling subtle spectral differences and large variations in spatial scale.
Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel cross-attention module.
To address these challenges, we propose MSCloudCAM, a cross-attention with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery. MSCloudCAM leverages complementary multi-scale context encoders and a newly designed cross-attention mechanism to dynamically fuse fine-grained spatial features with global contextual information. This design enables accurate pixel-wise classification across datasets with differing spectral responses, resolutions, and cloud characteristics.
Our key contributions are as follows:
  • We introduce a hybrid multi-scale context framework that leverages two complementary multi-scale context extractors. One emphasizes on fine-grained spatial details and the other on capturing broad global semantics. Thus it enables us to achieve comprehensive multi-scale feature utilization capability from a hierarchical vision backbone.
  • We design a novel convolution-based cross-attention module, i.e. CAM that leverages these complementary feature hierarchies through an adaptive formulation. This mechanism enables dynamic, scale-aware feature fusion, offering a more principled alternative to conventional stacking or concatenation.
  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [1] and Landsat-8 (L8Biome) [2], consistently outperforming state-of-the-art models while maintaining competitive computational efficiency.
II Related Work
II-A Cloud Segmentation in Remote Sensing
Cloud segmentation is essential for atmospheric correction, land cover mapping, and environmental monitoring. Early rule-based methods such as Fmask [3] relied on band thresholds and decision trees but frequently misclassified thin clouds, mixed pixels, and bright surfaces. Spectral index–based approaches (e.g., NDSI, NDVI, cloud probability) [4] provided improved robustness yet still struggled under strong spectral variability or semi-transparent clouds. Traditional machine learning classifiers (SVM, Random Forests, boosting) [5] incorporated spectral–textural cues but remained constrained by handcrafted features and limited context modeling, motivating the shift toward deep learning.
II-B Deep Learning Approaches
Deep learning has become the dominant strategy for cloud segmentation due to its ability to model spatial–spectral structure directly from multispectral inputs.
CNN-based architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP. Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
Transformer-based methods leverage self-attention for long-range dependency modeling, crucial for large or spatially diffuse clouds. SegFormer [16] uses a lightweight MiT backbone for global context, while Mask2Former [17] delivers strong generalization through masked attention. UPerNet-InternImage [18] combines hierarchical transformers with pyramid pooling to achieve state-of-the-art results in remote sensing segmentation.
Hybrid models merge CNNs for low-level edge and texture extraction with transformer blocks for global reasoning, including HRNet–Transformer variants and cross-attention designs [19].
Foundation-model–based segmentation adapts pretrained vision transformers such as DINOv2 [20] to remote sensing; however, effective use of multispectral data still requires spectral alignment and domain-specific fine-tuning.
II-C Gaps in Current Methods
Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization. Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention. Our MSCloudCAM addresses both challenges.
III Methodology
III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g. Swin Transformer [19] encoder with dual context extractor modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
Figure 2: Architecture of MSCloudCAM. A multispectral Swin Transformer encoder generates hierarchical features. ASPP and PSP modules enrich deep and intermediate representations. These are fused by convolutional cross-attention, refined via channel–spatial attention, and decoded by a multi-stage upsampling path with auxiliary supervision.
III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
$\{f_{1},f_{2},f_{3},f_{4}\}=\mathcal{E}(\mathbf{X}),\quad f_{i}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}},$ (1)
with channel dimensions $[96,192,384,768]$ and progressively reduced spatial sizes $(H_{i},W_{i})$. Shifted-window self-attention within Swin ensures that $f_{4}$ captures global dependencies, while $f_{1}$$f_{3}$ preserve finer detail.
III-C Multi-Scale Context Modules
To enhance representational richness, we employ two complementary multi context encoders that operate at different feature hierarchies.
As a part of dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
$x_{\text{ASPP}}$ $=\phi_{\text{ASPP}}(f_{4})$
$=\bigoplus_{r\in\{1,6,12,18\}}\sigma\!\big(\text{Conv}_{3\times 3}^{(r)}(f_{4})\big)\;\oplus\;\sigma\!\big(\text{GAP}(f_{4})\big),$ (2)
where $\oplus$ denotes channel-wise concatenation, $r$ specifies the dilation rate, GAP is global average pooling, and $\sigma$ represents the ReLU nonlinearity. This module injects large-scale semantic context while retaining spatial discrimination.
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing i.e. PSP [21] aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
$x_{\text{PSP}}=\phi_{\text{PSP}}(f_{3})=\bigoplus_{s\in\{1,2,3,6\}}\sigma\!\left(\text{Up}\big(\text{Pool}_{s}(f_{3})\big)\right),$ (3)
where $\text{Pool}_{s}$ denotes average pooling to an $s\times s$ grid, followed by projection and bilinear upsampling. This captures scene-level information at progressively coarser scales, which is essential for delineating fine structures such as thin clouds.
III-D Cross-Attention Fusion
The outputs from the two multi-scale context extractors are concatenated to form a joint feature representation:
$x_{\text{cat}}=[\,x_{\text{MC1}}\,\|\,x_{\text{MC2}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W},$ (4)
where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics. To effectively combine these complementary cues, we employ a convolutional cross-attention mechanism. Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure. Meanwhile, $x_{\text{MC2}}$ provides the keys and values, since its pooled multi-scale features offer stable, region-level guidance suitable for modulating finer details. This directional interaction allows each spatial location in $x_{\text{MC1}}$ to attend to the most relevant semantic context encoded by $x_{\text{MC2}}$.
Formally, the attention is computed as follows: the query $Q=W_{Q}*x_{\text{cat}}$, key $K=W_{K}*x_{\text{MC2}}$, and value $V=W_{V}*x_{\text{MC2}}$ are first obtained.
$\text{Attn}(Q,K,V)$ $=\operatorname*{softmax}\!\Big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\Big)V,$ (5)
$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}},$ (6)
where * denotes convolution and residual addition preserves the original spatial representation. This formulation enables fine-grained features to be adaptively enhanced by global contextual cues, producing a richer and more discriminative multi-scale representation.
III-E Bottleneck Projection
The fused representation $\tilde{x}$ is compressed and regularized through a three-layer bottleneck:
[…]
$))),\displaystyle=\sigma\!\Big(\text{Conv}_{1\times 1}\!\big(\sigma(\text{Conv}_{3\times 3}\!(\sigma(\text{Conv}_{1\times 1}(\tilde{x}$$)\big)\Big)$, (7)
which first reduces dimensionality, then captures local spatial dependencies, and finally projects features into a compact 512-channel embedding. This step reduces redundancy and prepares the representation for attention-based refinement.
III-F Combined Attention Refinement
The bottleneck output is recalibrated by a combined attention operator:
$z^{\prime}=\phi_{\text{CA}}(z)=\phi_{\text{ECA}}(z)\odot\phi_{\text{SA}}(z)+z,$ (8)
where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [22] and $\phi_{\text{SA}}$ denotes spatial attention [23, 24]. This joint formulation first reweights channels based on their global importance and then emphasizes spatially salient regions, allowing the model to focus on both discriminative spectral responses and structurally meaningful areas.
III-G Decoder with Deep Supervision
The refined representation $z^{\prime}$ is progressively upsampled through transposed convolutions:
$y_{i}=\sigma(\text{Deconv}_{i}(y_{i-1})),\quad i=1,\dots,4,$ (9)
with $y_{0}=z^{\prime}$. Two intermediate outputs, $\hat{Y}_{1}$ and $\hat{Y}_{2}$, are generated from $y_{1}$ and $y_{2}$, respectively, and resized to the input resolution $(H_{0},W_{0})$. The final prediction $\hat{Y}$ is obtained from $y_{4}$. The training objective jointly supervises all three outputs:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}(\hat{Y},Y)+\lambda_{1}\mathcal{L}(\hat{Y}_{1},Y)+\lambda_{2}\mathcal{L}(\hat{Y}_{2},Y),$ (10)
where $Y$ is the ground truth mask, and $(\lambda_{\text{final}},\lambda_{1},\lambda_{2})$ are supervision weights.
IV Experiments
TABLE I: Performance comparison on CloudSEN12 L2A (BOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [12] 85.42 42.76 80.80 53.62 65.65 92.14 59.90 89.38 69.81 77.81 94.83 53.53 90.27 63.45 75.52 86.82
CDNetv2 [10] 85.35 44.01 80.85 54.00 66.05 92.09 61.12 89.41 70.13 78.19 95.65 56.37 88.06 63.56 75.91 86.88
HRCloudNet [15] 87.24 44.18 82.78 59.20 68.35 93.19 61.28 90.58 74.37 79.85 95.41 51.00 93.17 69.83 77.35 88.35
U-Net-MobV2 88.74 83.87 54.13 60.15 71.72 94.03 91.23 70.24 75.12 82.65 93.45 95.55 93.71 95.43 94.53 89.07
DeepLabV3+-MobV2 88.01 83.70 50.36 61.78 70.96 93.62 91.12 66.98 76.37 82.03 92.93 95.46 93.83 95.45 94.42 88.84
DeepLabV3+-ResNet101 87.96 82.83 50.36 59.52 70.17 93.59 90.61 66.99 74.63 81.45 92.93 95.15 93.79 95.14 94.25 88.50
UPerNet-InternImage 87.79 82.22 40.31 63.96 68.57 93.30 89.78 51.63 77.19 77.98 93.08 95.55 93.48 95.50 94.40 88.81
Mask2Former-DINOv2 83.53 79.32 45.51 56.50 66.21 91.03 88.47 62.55 72.21 78.56 90.30 94.23 91.17 94.93 92.66 85.31
Mask2Former-Swin-T 88.91 84.35 50.72 63.25 71.81 94.13 91.51 67.30 77.49 82.61 93.61 95.77 93.33 95.30 94.50 89.01
SegFormer-MiT-B5 89.28 83.50 51.87 64.19 72.21 94.34 91.01 68.31 78.19 82.96 93.76 95.36 93.86 95.90 94.72 89.44
MSCloudCAM 89.28 84.30 55.53 63.38 73.12 94.33 91.48 71.40 77.58 83.70 93.76 95.69 94.20 95.74 94.85 89.69
IV-A Datasets
We evaluate MSCloud on two publicly available multispectral cloud segmentation datasets.
CloudSEN12 [1] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. L1C provides Top-of-Atmosphere (TOA) reflectance, while L2A provides Bottom-of-Atmosphere (BOA) surface reflectance. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
L8Biome [2] is derived from Landsat-8 CCA, containing 11 bands (30 m) with manual masks. The original five classes—fill (0), shadow (64), clear (128), thin cloud (192), and cloud (255)—are remapped to CloudSEN12’s taxonomy: clear $\rightarrow 0$, thick cloud $\rightarrow 1$, thin cloud $\rightarrow 2$, shadow $\rightarrow 3$. The fill class is ignored. Bands are normalized with the same $3000.0$ divisor for consistency.
TABLE II: Dataset split statistics for CloudSEN12 and L8Biome.
Dataset Train Validation Test
CloudSEN12 L1C 8,500 500 1,000
CloudSEN12 L2A 8,500 500 1,000
L8Biome 13,378 4,459 4,459
IV-B Loss Function
The training objective is a weighted sum of the losses from the final prediction and the two auxiliary predictions:
$\mathcal{L}=\lambda_{\text{final}}\mathcal{L}_{\text{final}}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2}.$ (11)
We set $\lambda_{\text{final}}=1.0$ and $\lambda_{1}=\lambda_{2}=0.4$, ensuring the final prediction dominates while auxiliary supervision stabilizes early-stage training.
IV-C Training Protocol
Our model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-4}$. The batch size is set to 8, and we train for 100 epochs.
IV-D Evaluation Metrics
$\mathrm{IoU}_{i}$ $=\frac{\mathrm{TP}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (12)
$\mathrm{F1}_{i}$ $=\frac{2\,\mathrm{TP}_{i}}{2\,\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}},$ (13)
$\mathrm{Acc}_{i}$ $=\frac{\mathrm{TP}_{i}+\mathrm{TN}_{i}}{\mathrm{TP}_{i}+\mathrm{FP}_{i}+\mathrm{FN}_{i}+\mathrm{TN}_{i}}.$ (14)
Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
V Results
V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIIIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L1C (TOA).
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
DBNet [12] 85.20 43.02 81.40 52.44 65.52 92.01 60.16 89.75 68.80 77.68 95.56 53.15 89.40 62.47 75.15 86.83
CDNetv2 [10] 84.98 43.37 80.67 53.40 65.60 91.88 60.50 89.30 69.62 77.83 94.80 53.13 89.32 65.34 75.65 86.68
HRCloudNet [15] 86.01 45.88 83.54 57.61 68.26 92.48 62.90 91.03 73.10 79.88 94.49 59.47 91.39 67.19 78.13 87.86
U-Net-MobV2 90.82 85.98 50.54 67.21 73.64 95.19 92.46 67.14 80.39 83.80 94.42 96.11 95.05 96.62 95.55 91.10
DeepLabV3+-MobV2 90.13 85.75 48.97 66.29 72.79 94.81 92.33 65.75 79.73 83.15 94.02 96.04 94.53 96.51 95.28 90.55
DeepLabV3+-ResNet101 90.65 85.66 50.60 66.36 73.32 95.10 92.28 67.20 79.78 83.59 94.38 95.94 94.94 96.47 95.43 90.87
UPerNet-InternImage 90.85 85.28 36.70 68.87 70.42 95.09 91.82 47.94 80.98 78.96 94.50 96.43 95.24 96.85 95.75 91.51
Mask2Former-DINOv2 88.50 82.10 43.61 63.62 69.46 93.90 90.17 60.73 77.76 80.64 92.94 94.84 94.41 96.11 94.57 89.15
Mask2Former-Swin-T 90.75 86.70 49.90 69.53 74.22 95.15 92.88 66.58 82.02 84.16 94.46 96.32 94.34 96.93 95.51 91.03
SegFormer-MiT-B5 91.47 86.70 49.77 70.24 74.54 95.55 92.88 66.46 82.52 84.35 94.84 96.27 95.25 96.92 95.82 91.64
MSCloudCAM 91.64 87.24 52.58 70.63 75.52 95.64 93.19 68.92 82.78 85.13 94.93 96.42 95.58 97.04 95.99 91.99
TABLE IV: Performance comparison on L8Biome.
Model IoU (%) F1 (%) Accuracy (%) aAcc
Clear Thick Thin Shadow mIoU Clear Thick Thin Shadow mF1 Clear Thick Thin Shadow mAcc (%)
CDNetv2 [10] 73.81 24.85 75.85 00.00 43.63 84.93 39.81 86.27 00.00 52.75 90.18 31.95 89.81 00.00 52.98 78.56
HRCloudNet [15] 72.41 30.89 70.76 00.00 43.51 83.99 47.20 82.88 00.00 53.52 85.50 43.70 85.88 00.00 53.77 77.04
DBNet [12] 80.19 38.18 82.37 04.90 51.41 89.00 55.25 90.34 09.34 60.99 96.51 49.34 87.89 05.03 59.70 83.62
U-Net-MobV2 88.98 78.59 39.97 6.37 53.48 94.17 88.01 57.11 11.98 62.82 94.26 91.50 87.55 98.55 92.97 85.93
DeepLabV3+-MobV2 87.25 79.21 33.37 2.54 50.59 93.19 88.40 50.05 4.96 59.15 93.32 91.25 87.50 98.54 92.65 85.31
DeepLabV3+-ResNet101 87.56 74.19 34.24 4.88 50.22 93.37 85.18 51.01 9.31 59.72 93.47 89.56 85.51 98.56 91.78 83.55
UPerNet-InternImage 54.68 46.97 18.65 1.56 30.46 60.19 53.40 26.72 2.55 35.72 91.00 90.58 83.73 98.35 90.92 81.83
Mask2Former-DINOv2 69.67 70.75 34.67 13.13 47.06 82.12 82.87 51.49 23.22 59.93 84.03 86.86 84.39 98.65 88.48 76.96
Mask2Former-Swin-T 84.36 78.64 37.50 13.54 53.51 91.52 88.04 54.55 23.85 64.49 91.74 91.62 86.01 98.62 92.00 84.00
SegFormer-MiT-B5 88.62 74.38 39.52 15.27 54.45 93.97 85.31 56.65 26.50 65.61 94.04 90.12 86.05 98.70 92.23 84.45
MSCloudCAM 88.27 82.23 40.19 21.48 58.04 93.77 90.25 57.34 35.36 69.18 93.84 92.91 88.74 98.70 93.55 87.09
V-B Qualitative Results
To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results. Each row shows a different dataset, and each column a different method.
Figure 3: Qualitative cloud segmentation results. Each row corresponds to a dataset. The proposed method produces sharper delineation of thin clouds and shadows compared to other approaches. Class–color mapping: Class 1 (Clear sky – light blue), Class 2 (Thick cloud – white), Class 3 (Thin cloud – light gray), Class 4 (Cloud shadow – dark gray).
V-C Ablation Study
To evaluate the role of each multicontext component, we conduct ablation experiments by selectively replacing or removing modules. Replacing ASPP with PSP enhances global context but loses fine detail, while using ASPP instead of PSP improves multi-scale local features but sacrifices global scene understanding.
Using PSP and ASPP together provides complementary global–local context and outperforms either module alone. Removing the Cross-Attention Module and using simple concatenation weakens the model’s ability to emphasize informative regions. The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy.
TABLE V: Ablation study evaluating dual multi context extractor configurations and our cross-attention module (CAM) on CloudSEN12 L2A (BOA).
Dual Extractor Fusion Type mIoU mDice mAcc aAcc
PSP + PSP CAM 70.26 81.61 92.05 86.41
ASPP + ASPP CAM 70.09 81.25 91.93 86.18
PSP + ASPP Concat 71.48 82.44 93.52 88.02
PSP + ASPP CAM 73.12 83.70 94.85 89.69
V-D Model Complexity Analysis
Table VI summarizes the computational complexity of all evaluated models in terms of FLOPs, parameter counts, and inference efficiency. In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance. All experiments were conducted on a single NVIDIA A100 40GB GPU using PyTorch 2.3 and CUDA 12.1, with a batch size of 8 and an input resolution of 512×512. All models were evaluated in FP32 precision. Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM module only adds 1.7M trainable parameters to the model.
TABLE VI: Model complexity comparison in terms of FLOPs (G), parameters (M), inference time (ms/image), and TP as throughput (FPS).
[…]
U-Net-MobileNetV2 13.6 6.6 11.5 87.0
DeepLabV3+-MobileNetV2 6.1 4.3 14.8 67.5
DeepLabV3+-ResNet101 56.1 45.6 38.5 26.0
UPerNet-InternImageDCNv4 193.5 110.8 52.0 19.2
Mask2Former-Swin-T 73.09 47.4 33.6 29.7
Mask2Former-DINOv2-B14 85.6 90.6 44.0 22.7
SegFormer-MiT-B5 55.4 45.6 28.7 34.8
MSCloudCAM 38.98 47.44 15.9 62.7
VI Code and Model Availability
The full implementation of our proposed MS-CloudCAM framework is publicly available. The source code, training scripts, and experimental configurations are hosted on GitHub: https://github.com/mazid-rafee/ms-cloudcam. Additionally, pretrained model checkpoints are released on Hugging Face: https://huggingface.co/mazid-rafee/MS-CloudCAM/tree/main.
VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating hierarchical vision backbone with dual multi-scale context extractors and emlpoying our novel convolution based cross attention module, our method achieves superior segmentation accuracy over state-of-art-models on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.
Acknowledgment
This material is based in part upon work supported by the National Science Foundation under Grants CNS-2018611 and CNS-1920182.